$2000 loan

How much interest do you end up paying if you make 45.214 payments of $60?

How much interest do you end up paying if you make 23.6285 payments of $100 instead?
Answer #1

Loan = 2000

For 45.214 payment of $60 each,

Total payment = 45.214*60 = $2712.84

Interest payment = $2712.84-2000 = $712.84

For 23.6585 payments of 100 each

Total payment = 23.6585*100 = $2365.85

Interest payment = $2365.85-2000 = $365.85